OpenStudy (anonymous):

a=6 b=10 c=? what does c equal???

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

use Pythagorean thereom (retriceassuming this is a triangle) a^2+b^2=c^2

OpenStudy (anonymous):

is that 6squared?

OpenStudy (anonymous):

@~Jane_Doe~

OpenStudy (anonymous):

you would set up the equation as 6^2+10^2=c^2 solve for c....yes ^2= 2 squared